# Filtering Spin and Orbital Moment in Centrosymmetric Systems

### v2

@@ Appendix B Solving the transmission problem
i.e., the linear combination of transmitted and reflected plane waves with the electronic orbital and spin character contained in the vector $\left|\phi_{j}^{A}\right\rangle$. Here, the generic electron wavenumber $k_{j}^{A}$ is expressed in units of $k_{0}$. In our case, we put $\left|\Psi_{in}^{\text{I}}\left(z\right)\right\rangle=\sum_{j=1}^{6}\alpha_{j}^{\text{I}}e^{ik_{j}^{\text{I}}z}\left|\phi_{j}^{\text{I}}\right\rangle$; in addition, assuming there is only a transmitted wavefunction in region III, we pose $\beta_{j=1,\ldots,6}^{\text{III}}=0$. The state has to be continuous in space, as well as its first derivative when $\hat{\mathcal{A}}^{\text{I}}=\hat{\mathcal{A}}^{\text{II}}=\hat{\mathcal{A}}^{\text{III}}$ and there are no singularities in space. Therefore, the following conditions have to be fulfilled simultaneously:
@@ Appendix C Transmission Matrix and representative physical cases
^$\hat{\mathcal{H}}\left(z\right)=\begin{cases}\hat{p}_{z}^{2}\text{$\mathcal{I}$}&z<-d/2\\ \hat{p}_{z}^{2}\left(\text{$\mathcal{I}$}+g_{0x}\hat{\chi}_{x}\right)&-d/2\leq z\leq d/2\\ \hat{p}_{z}^{2}\text{$\mathcal{I}$}&z>d/2\end{cases},$ (27)
@@
^$\hat{\mathcal{H}}\left(z\right)=\begin{cases}\hat{p}_{z}^{2}\mathcal{I}+V\left(z\right)&z<-d/2\\ \hat{p}_{z}^{2}\mathcal{I}+g_{0x}\hat{\chi}_{x}+V\left(z\right)&-d/2\leq z\leq d/2\\ \hat{p}_{z}^{2}\mathcal{I}+V\left(z\right)&z>d/2\end{cases},$ (43)

### v3

@@ Appendix B Solving the transmission problem
i.e., the linear combination of transmitted and reflected plane waves with the electronic orbital and spin character contained in the vector $\left|\phi_{j}^{A}\right\rangle$. Here, the generic electron wavenumber $k_{j}^{A}$ is expressed in units of $k_{0}$. In our case, we put $\left|\Psi_{in}^{\text{I}}\left(z\right)\right\rangle=\sum_{j=1}^{6}\alpha_{j}^{\text{I}}e^{ik_{j}^{\text{I}}z}\left|\phi_{j}^{\text{I}}\right\rangle$; in addition, assuming there is only a transmitted wavefunction in region III, we pose $\beta_{j=1,\ldots,6}^{\text{III}}=0$. The state has to be continuous in space, as well as its first derivative when $\hat{\mathcal{A}}^{\text{I}}=\hat{\mathcal{A}}^{\text{II}}=\hat{\mathcal{A}}^{\text{III}}$ and there are no singularities in space ($V_{1}=V_{2}=0$). Therefore, the following conditions have to be fulfilled simultaneously:
@@ Appendix C Transmission Matrix and representative physical cases
^$\hat{\mathcal{H}}\left(z\right)=\begin{cases}\hat{p}_{z}^{2}\text{$\mathcal{I}$}&z<-d/2\\ \hat{p}_{z}^{2}\text{$\mathcal{I}$}+g_{0x}\hat{\chi}_{x}&-d/2\leq z\leq d/2\\ \hat{p}_{z}^{2}\text{$\mathcal{I}$}&z>d/2\end{cases},$ (41)
@@
^$\hat{\mathcal{H}}\left(z\right)=\begin{cases}\hat{p}_{z}^{2}\mathcal{I}&z<-d/2\\ \hat{p}_{z}^{2}\mathcal{I}+g_{0x}\hat{\chi}_{x}+V\left(z\right)\mathcal{I}&-d/2\leq z\leq d/2\\ \hat{p}_{z}^{2}\mathcal{I}&z>d/2\end{cases},$ (57)